JuicyTomato:

If ƒ (x ) = 3x2 - x, find ƒ (-2).

ninjabeeler1234:

f(x)=3x*2 so f(-2)=-12 but if you meant f(x)=3*2. then f(-2)=6.

jhonyy9:

so i think this function wan be f(x ) = 3x^2 -x so to calcule f(-2) just you need substitute the -2 in place of x inside this equation and make the calcule hope helped understandably easy